OpenStudy (anonymous):

Ok, here's another, don't kill me you guys:( x/4-x/5=1, I think the answer is 10 but it isn't checking out when i replace the x's with 10:( Help??

OpenStudy (inkyvoyd):

x/4-x/5=1 5x-4x=20 x=20

OpenStudy (inkyvoyd):

(step two: multiply both sides by 20)

OpenStudy (anonymous):

I did get that far so let me try again;)

OpenStudy (anonymous):

you can just change x/5 to 5x?

OpenStudy (inkyvoyd):

No, you multiply both sides by 20,

OpenStudy (inkyvoyd):

20(x/4)-20(x/5)=20(1)

OpenStudy (inkyvoyd):

5x-4x=20

OpenStudy (anonymous):

I'm lost, thank you anyway.

OpenStudy (inkyvoyd):

Multiply each of the terms by 20

OpenStudy (anonymous):

Ok, I'll work it out, thanks

OpenStudy (anonymous):

Oh I get it, ok, thank you much, Denise

OpenStudy (inkyvoyd):

:)